Purchase Government bond which currently trades at TZS 850 with 10% annual coupon rate. The bond was issued on 12nd January 2pe you Required o What is the price of Bond. (2) Profitability index (3) And other Glaulation that is relevant
Answer #1

2. Profitability index= Present value of future cash flows-initial investment

From step 1, we have found the bond price to be 971.27 TZSBond pice= Cx T-(I+t) Cltr) F = 850. C= 101. = 850x10:1- =85 r/YTM=8.5% 85x81-(1+ 0:085)- 850 %3D n= 20 20 Price (1+0.085)²

therefore PI=971.27-850=1.14

profitability index which is <= 1 is considered acceptable